(3 points) How many integer solutions of + +33 +34 = 30 are there with I, > 0, 2 <5, 12 <11, 13 < 13, 14 < 23?
Answer #1

07 A.ccording to question we have to find how many is possible for integer solution (A) Hq t U + Ug + My =30 where Nizo, 4, 5Similar m formula Mit Ut isk Yui =n + kk = iil and Mizo then this is a case of n identical number ( like I, I, upto ..n.) whpop let Change => 0_5-24 -) 5-34120 ilet 5-2, =P (plug this pealue = p zo 5-PE Exi Aimilarly H2 <1 = 117.HQ XO Let 11- = 0 ag- 5-P + 11-9 + 13-11+ 23-5 Mit Ng + H3 + Hy = 30 09 = 30 52- (P+9+9++) = 30 =) Ptq + rtt 52-30 22 Ly 8. P+9+977 + = again her© Doy n= 221 k=u 22+ 4-1 total number of integer solution Curl (22+4-1)! (22)! (4-1)! こ 25! (22)! 3! 25x24x23x022) (22)! 3X2X

May be helpful for you....!!! And try to understand the concept.

Please upvote me....!!